# Holographic thermodynamics of a five-dimensional neutral Gauss-Bonnet AdS black hole

### v1

@@ Abstract
Motivated by the recent progress on the holographic dual of the extended thermodynamics for black holes in anti-de Sitter (AdS) space, we investigate the holographic thermodynamics for the five-dimensional neutral Gauss-Bonnet AdS black hole in the context of the anti–de Sitter/conformal field theory (AdS/CFT) correspondence. Through the extended bulk thermodynamics for the five-dimensional Gauss-Bonnet AdS black hole, we derive the first law of the CFT thermodynamics which is obtained by directly translating the arbitrary conformal factors in the dual CFT. In addition to the newly defined chemical potential $\mu$ conjugating to the central charge $C$, we obtain other pairs of thermodynamics for the CFT, such as the temperature $\tilde{T}$ and the entropy $S$, the Gauss-Bonnet coupling constant $\tilde{\alpha}$ and its conjugate variable $\tilde{\mathcal{A}}$, the pressure $\mathcal{P}$ and its conjugate volume $\mathcal{V}$. In the fixed $C$, $\mathcal{V}$ and $\tilde{\alpha}$ canonical ensemble, we obtain the canonical description of the CFT thermodynamics and observe the standard swallowtail behavior in the Helmholtz free energy vs the temperature plot. The self-intersection point of the Helmholtz free energy indicates the phase transition between the high and low entropy states of the CFT. By using Maxwell’s equal area law, we get the critical point and coexistence curve for the high and low entropy phases of the CFT. Besides, we get the critical exponents for the CFT, and find that the critical point and critical exponents associated with the $\tilde{T}-S$ criticality of the CFT are the same as those of the five-dimensional Gauss-Bonnet AdS black hole.
Keywords:
black hole thermodynamics, holographic, AdS/CFT correspondence.
1 Introduction
@@
Black hole thermodynamics in asymptotically anti-de Sitter (AdS) spacetime has advanced to the forefront of current research activities in theoretical high energy physics. One of the intriguing aspects of theoretical physics is duality which has revolutionized perspectives of physics at a fundamental level. Dualities relate two different facets of nature and thus they lead to an enhanced understanding of physics. An implication of these dualities is that, while the mathematical structures of the two theories may be identical, there may exist differences from a physical standpoint. One of the interesting dualities is the anti-de Sitter/conformal field theory (AdS/CFT) correspondence Maldacena:1997re ; Gubser:1998bc ; Witten:1998qj which connects the CFT on a flat spacetime to string theory. This is interesting because string theory is a theoretical framework for quantum gravity, whereas quantum field theory is not.
@@
Now, even though these two theories are different, the AdS/CFT correspondence assumes the existence of a correspondence between the two theories and this is one of the most attractive features of this duality, which pushes research for its deeper understanding and possible applications. A renowned example of a phase transition in black hole spacetimes is the radiation/black hole first-order phase transition of Hawking and Page, observed for Schwarzschild-AdS black holes immersed in a bath of radiation HaPa83 . Such a phenomenon has a dual interpretation for a boundary quantum field theory via the AdS/CFT correspondence and is related to a confinement/deconfinement phase transition in the dual quark gluon plasma Witt98 ; Witten:1998qj . For charged black hole in AdS spacetime one likewise observes a small/large black hole first-order phase transition reminiscent of the liquid/gas phase transition of the Van der Waals fluid CEJM99a ; CEJM99b . But when the cosmological constant is interpreted as the thermodynamic pressure and the black hole mass is interpreted as the thermodynamic enthalpy KaRT09 ; Dola11 , the phase behavior of the charged AdS black hole is exactly the same with that of the Van der Waals fluid and they have the same critical exponents KuMa12 . This version of the black hole thermodynamics in the presence of negative cosmological constant is interpreted as the extended phase space thermodynamics or generically speaking, the black hole chemistry KuMT17 ; Mann:2016trh . The phase structures in the extended phase space is much richer, including reentrant phase transition Gunasekaran:2012dq ; Frassino:2014pha , superfluid-like phase transition Hennigar:2016xwd , triple points Altamirano:2013uqa ; Wei:2014hba ; Altamirano:2013ane etc. Recent advancements have yielded numerous successful interpretations of the origins of the extended phase space thermodynamics. Notable among these are the extended Iyer-Wald formalism XiTL24 , the coupling constant of gauge field perspective HaTe24 . Additionally, a mechanism for the higher-dimensional origin of a dynamical cosmological constant was proposed fpsv23 .
@@
Despite significant advancements have been achieved in the field of thermodynamics for black holes in AdS spacetime, the holographic interpretation of black hole chemistry has remained an enigma Mann24 . From the viewpoint of the AdS/CFT correspondence, the negative cosmological constant $\Lambda$ sets the asymptotic structure of the bulk spacetime and is related to the number of colours $N$ in the dual gauge field theory Mald97 ; KaRo15
@@
where $L$ is the AdS radius and the numerical factor $k$ depends on the details of the particular holographic system. Hence, variation of the negative cosmological constant $\Lambda$ corresponds to changing the number of colors $N$ and hence varying the dual gauge field theories in the boundary Johnson:2014yja ; CKMV22 ; CoKM21 .
Recently there have been considerable progress in developing an exact dictionary between the black hole chemistry and the thermodynamics of the dual gauge field theory. Considering the viewpoint that Newton’s constant is treated as a variable CoKM21 , and extended to the more general CFT metric with boundary radius different from the AdS radius, Visser argued that holographic thermodynamics requires a chemical potential for color and obtained the thermodynamic Euler equation for high energy states of large-$N$ gauge theories Viss22 . Systematic works have been done based on this viewpoint Zheng:2024glr ; Zhang:2023uay ; Sadeghi:2023tuj ; Ali:2023jox ; Punia:2023ilo . Building on Visser’s viewpoint that the radius of the CFT metric need not coincide with the AdS radius, and allowing the conformal factor to vary while keeping Newton’s constant fixed, thereby resulting in the boundary volume $\mathcal{V}$ and the central charge $C$ being completely independent, Ahmed et al. constructed an exact dictionary between the thermodynamics of the charged AdS black hole and the dual conformal field theory ACKMV23a . Based on this perspective, a lot of work has been done for the holographic thermodynamics for AdS black holes, including the charged AdS black hole Ladghami:2024wkv , Kerr AdS black hole GoJZ23 ; ACKMV23 , Kerr-Newman AdS black hole Baruah:2024yzw . Inspired by this viewpoint, we investigate the holographic CFT thermodynamics for the five-dimensional neutral Gauss-Bonnet AdS black hole. Different from the recent work of Cui et al. CuFa24 , where the Newton’s constant is regarded as a dynamical variable, we take the viewpoint of Ahmed et al. in which the Newton’s constant is treated as a fixed quantity but the conformal factor is supposed to vary.
@@
The outline of the paper is as follows. In Sec. 2, we derive the AdS/CFT dictionary and the laws for the CFT residing in the boundary dual to the five-dimensional Gauss-Bonnet AdS black hole. In Sec. 3, we investigate the phase behavior and criticality of the CFT. Finally, in the last section we discuss the results and conclude them.
2 Holographic thermodynamics of five-dimensional Gauss-Bonnet AdS black hole
In this section, we tentatively establish a connection between the extended thermodynamics of the five-dimensional neutral Gauss-Bonnet AdS black hole and the corresponding extended thermodynamics of dual CFT. By exploring the relationship between these two realms, we aim to shed light on the profound similarities and underlying principles that govern their thermodynamic properties.
2.1 Extended bulk thermodynamics
In this subsection, we briefly review the main points of the extended black hole thermodynamics for the five-dimensional neutral Gauss-Bonnet AdS black hole. A $d$-dimensional neutral Gauss-Bonnet AdS black hole is a solution for the gravitational field equation derived by the action BoDe85
@@ 2.1 Extended bulk thermodynamics
Within the context of the low energy effective action formalism in heterotic string theory, the Gauss–Bonnet term is the leading order quantum correction to gravity GrSl87 . The Gauss–Bonnet coupling constant is related to the string scale and can be identified with the inverse of the string tension ChDu02 ; Wei:2012ui ; YWCYW20 . Hence, we only consider positive Gauss-Bonnet coupling constant in this paper.
The theory admits an exact neutral black hole solution in five dimensions. The metric for the five-dimensional neutral Gauss-Bonnet AdS black hole is Cai02
$ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+r^{2}[d\theta^{2}+\sin^{2}\theta(d\phi^{2}+\sin^{2}\phi d\psi^{2})]$ (5)
with the metric function given by
@@
From the viewpoint of the extended black hole thermodynamics, the cosmological constant $\Lambda$ is interpreted as the thermodynamic pressure and the black hole mass is regarded as the thermodynamic enthalpy $H$ KaRT09 . The thermodynamic pressure is
$P=-\frac{\Lambda}{8\pi G_{\text{N}}}=\frac{3}{4\pi G_{\text{N}}}\frac{1}{L^{2}},$ (9)
and its conjugate thermodynamic volume is
@@
The Gauss-Bonnet coupling constant $\alpha$ is a thermodynamic variable, with the corresponding conjugate thermodynamic variable given by
@@
Evidently, the presence of the Gauss-Bonnet correction to Einstein’s theory of gravity leads to a notable revision of the Bekenstein-Hawking area law when considering the black hole’s temperature as its surface gravity, as emphasized in Ref. KuLi23 .
The first law of black hole thermodynamics and the Smarr relation for the five-dimensional Gauss-Bonnet AdS black hole are CCLY13
@@
Using the first law of the bulk thermodynamics, we proceed to derive the boundary thermodynamics in the next subsection. This derivation establishes a tentative relationship between the thermodynamic properties of the bulk system and its boundary counterpart, allowing for a comprehensive analysis of their interconnected behavior. The insights gained from this derivation might enhance our understanding on the overall thermodynamic framework.
2.2 Extended boundary thermodynamics
From the perspective of the AdS/CFT correspondence, the dual CFT lives on the conformal boundary of the asymptotically AdS spacetime. According to the recent development of the holographic chemistry Mann24 , the boundary metric of the CFT dual to the five-dimensional Gauss-Bonnet AdS spacetime is Gubser:1998bc ; Karc15 :
@@ 2.2 Extended boundary thermodynamics
and $\omega=R/L$ is the dimensionless conformal factor that is free to vary Viss22 ; ACKMV23a ; ACKMV23 . It reflects the conformal symmetry of the boundary theory. $R$ is the radius of the boundary.
In the gauge/gravity duality, the central charge of the CFT is related to the AdS radius $L$ and the gravitational coupling constant $G_{\text{N}}$ in the bulk theory. For Einstein gravity the central charge usually reads CKMV22
$C=\frac{\Omega_{d-1}L^{d-1}}{16\pi G_{\text{N}}},$ (17)
where $\Omega_{d-1}$ is the volume of a $(d-1)$-dimensional unit sphere. Since the precise coefficient of the central charge is irrelevant for our purpose and only the combination $dC/C$ appears in the first law of the CFT, the dimensional dependent coefficient is normalized such that it agrees with the coefficient $A$ of the Euler density in the trace anomaly CKMV22 ; Myers:2010tj ; Chen:2023pgs .
From the first law of black hole thermodynamics and the Smarr relation for the five-dimensional Gauss-Bonnet AdS black hole, we can express the first law in terms of the boundary thermodynamics
$\begin{split}\delta\left(\frac{M}{\omega}\right)=\left(\frac{T}{\omega}\right)\delta S+\frac{M-TS-\mathcal{A}\alpha}{\omega}\frac{\delta\left(L^{3}/G_{\text{N}}\right)}{\left(L^{3}/G_{\text{N}}\right)}+\frac{\sqrt{G_{\text{N}}}\mathcal{A}}{\omega L}\delta\left(\frac{\alpha L}{\sqrt{G_{\text{N}}}}\right)-\left(\frac{\tilde{E}}{3\mathcal{V}}\right)\delta\mathcal{V},\end{split}$ (18)
where $\mathcal{V}=2\pi^{2}R^{3}$ is the spatial volume of the boundary on which the CFT resides. In the derivation of the first law in terms of the boundary thermodynamics, the Newtown constant $G_{\text{N}}$ is regarded as a fixed constant, but the conformal factor $\omega$ is free to vary. Employing the following holographic dictionary:
$\begin{split}\tilde{E}&=\frac{M}{\omega},\qquad\qquad\qquad\qquad\tilde{T}=\frac{T}{\omega},\qquad\qquad\qquad S=S,\\ \tilde{\mathcal{A}}&=\frac{\sqrt{G_{\text{N}}}\mathcal{A}}{\omega L},\qquad\qquad\qquad\tilde{\mathcal{\alpha}}=\frac{\alpha L}{\sqrt{G_{\text{N}}}},\qquad\qquad\mathcal{P}=\frac{\tilde{E}}{3\mathcal{V}},\\ C&=\frac{\pi L^{3}}{8G_{\text{N}}},\qquad\qquad\qquad\mu=\frac{1}{C}\left(\tilde{E}-\tilde{T}S-\tilde{\mathcal{A}}\tilde{\mathcal{\alpha}}\right),\end{split}$ (19)
@@
    $\tilde{\alpha}=\sqrt{\frac{8CL^{3}}{\pi}}y,\qquad\tilde{\mathcal{A}}=-\frac{1}{R}\sqrt{\frac{\pi C}{8L^{3}}}\frac{24x^{4}+9x^{2}-6y}{x^{2}+2y}.$ (28)
@@
3 Thermodynamic ensembles in the dual CFT
The extended thermodynamics and phase transition for CFT states that are dual to the singly-spinning asymptotically AdS black hole were investigated by Ahmed et al., and various phase transitions and critical behavior in the CFT were uncovered ACKMV23 . Gong et al. found that the critical exponents for the CFT states in the vicinity of the critical point are the same with that of the Kerr-Newman AdS black hole GoJZ23 . In this section, we investigate the phase behavior and criticality exhibited by the canonical ensemble within the CFT framework.
3.1 Canonical ensemble: swallowtail behavior and critical point
The canonical ensemble for five-dimensional Gauss-Bonnet AdS black hole has been well study in the framework of extended black hole thermodynamics, i.e. black hole chemistry CCLY13 ; WeLi20 . In these studies it was found that the five-dimensional Gauss-Bonnet AdS black hole displays the swallowtail behavior, which indicates the small/large black hole phase transition reminiscent of the Van der Waals like phase transition. Besides, it was shown that the critical point for the five-dimensional Gauss-Bonnet AdS black hole phase transition is CCLY13

### v2

@@ Abstract
Motivated by the recent progress on the holographic dual of the extended thermodynamics for black holes in anti-de Sitter (AdS) space, we investigate the holographic thermodynamics for the five-dimensional neutral Gauss-Bonnet AdS black hole in the context of the anti-de Sitter/conformal field theory (AdS/CFT) correspondence. Through the extended bulk thermodynamics for the five-dimensional Gauss-Bonnet AdS black hole, we derive the first law of the CFT thermodynamics which is obtained by directly translating the arbitrary conformal factors in the dual CFT. In addition to the newly defined chemical potential $\mu$ conjugating to the central charge $C$, we obtain other pairs of thermodynamics for the CFT, such as the temperature $\tilde{T}$ and the entropy $S$, the Gauss-Bonnet coupling constant $\tilde{\alpha}$ and its conjugate variable $\tilde{\mathcal{A}}$, the pressure $\mathcal{P}$ and its conjugate volume $\mathcal{V}$. In the fixed $C$, $\mathcal{V}$ and $\tilde{\alpha}$ canonical ensemble, we obtain the canonical description of the CFT thermodynamics and observe the standard swallowtail behavior in the Helmholtz free energy vs the temperature plot. The self-intersection point of the Helmholtz free energy indicates the phase transition between the high and low entropy states of the CFT. By using Maxwell’s equal area law, we get the critical point and coexistence curve for the high and low entropy phases of the CFT. Besides, we get the critical exponents for the CFT, and find that the critical point and critical exponents associated with the $\tilde{T}-S$ criticality of the CFT are the same as those of the five-dimensional Gauss-Bonnet AdS black hole.
I Introduction
@@
Black hole thermodynamics in asymptotically anti-de Sitter (AdS) spacetime has advanced to the forefront of current research activities in theoretical high energy physics. One of the intriguing aspects of theoretical physics is duality which has revolutionized perspectives of physics at a fundamental level. Dualities relate two different facets of nature and thus they lead to an enhanced understanding of physics. An implication of these dualities is that, while the mathematical structures of the two theories may be identical, there may exist differences from a physical standpoint. One of the interesting dualities is the anti-de Sitter/conformal field theory (AdS/CFT) correspondence [1, 2, 3] which connects the CFT on a flat spacetime to string theory. This is interesting because string theory is a theoretical framework for quantum gravity, whereas quantum field theory is not.
@@
Now, even though these two theories are different, the AdS/CFT correspondence assumes the existence of a correspondence between the two theories and this is one of the most attractive features of this duality, which pushes research for its deeper understanding and possible applications. A renowned example of a phase transition in black hole spacetimes is the radiation/black hole first-order phase transition of Hawking and Page, observed for Schwarzschild-AdS black holes immersed in a bath of radiation [4]. Such a phenomenon has a dual interpretation for a boundary quantum field theory via the AdS/CFT correspondence and is related to a confinement/deconfinement phase transition in the dual quark gluon plasma [5, 3]. Further investigation of Hawking-Page phase transition for higher dimensional Schwarzshild-like AdS black hole in higher derivative gravity from the AdS/CFT correspondence perspective can be found in Refs. [6, 7, 8]. For charged black hole in AdS spacetime one likewise observes a small/large black hole first-order phase transition reminiscent of the liquid/gas phase transition of the Van der Waals fluid [9, 10]. But when the cosmological constant is interpreted as the thermodynamic pressure and the black hole mass is interpreted as the thermodynamic enthalpy [11, 12], the phase behavior of the charged AdS black hole is exactly the same with that of the Van der Waals fluid and they have the same critical exponents [13]. This version of the black hole thermodynamics in the presence of negative cosmological constant is interpreted as the extended phase space thermodynamics or generically speaking, the black hole chemistry [14, 15]. The phase structures in the extended phase space is much richer, including reentrant phase transition [16, 17], superfluid-like phase transition [18], triple points [19, 20] etc. Recent advancements have yielded numerous successful interpretations of the origins of the extended phase space thermodynamics. Notable among these are the extended Iyer-Wald formalism [21], the coupling constant of gauge field perspective [22]. Additionally, a mechanism for the higher-dimensional origin of a dynamical cosmological constant was proposed [23].
@@
Despite significant advancements have been achieved in the field of thermodynamics for black holes in AdS spacetime, the holographic interpretation of black hole chemistry has remained an enigma [24]. From the viewpoint of the AdS/CFT correspondence, the negative cosmological constant $\Lambda$ sets the asymptotic structure of the bulk spacetime and is related to the number of colours $N$ in the dual gauge field theory [25, 26]
@@
where $L$ is the AdS radius and the numerical factor $k$ depends on the details of the particular holographic system. Hence, variation of the negative cosmological constant $\Lambda$ corresponds to changing the number of colors $N$ and hence varying the dual gauge field theories in the boundary [27, 28, 29].
Recently there have been considerable progress in developing an exact dictionary between the black hole chemistry and the thermodynamics of the dual gauge field theory. Considering the viewpoint that Newton’s constant is treated as a variable [29], and extended to the more general CFT metric with boundary radius different from the AdS radius, Visser argued that holographic thermodynamics requires a chemical potential for color and obtained the thermodynamic Euler equation for high energy states of large-$N$ gauge theories [30]. Systematic works have been done based on this viewpoint [31, 32, 33, 34, 35]. Building on Visser’s viewpoint that the radius of the CFT metric need not coincide with the AdS radius, and allowing the conformal factor to vary while keeping Newton’s constant fixed, thereby resulting in the boundary volume $\mathcal{V}$ and the central charge $C$ being completely independent, Ahmed et al. constructed an exact dictionary between the thermodynamics of the charged AdS black hole and the dual conformal field theory [36]. Based on this perspective, a lot of work has been done for the holographic thermodynamics for AdS black holes, including the charged AdS black hole [37], Kerr AdS black hole [38, 39], Kerr-Newman AdS black hole [40]. Building on this perspective, we extend the study of the holographic dictionary to five-dimensional Gauss-Bonnet gravity, with a focus on the holographic thermodynamics and a comparison to the thermodynamics of the five-dimensional neutral Gauss-Bonnet AdS black hole. In our analysis, we treat Newton’s constant as a fixed parameter, while allowing the conformal factor to vary.
@@
The outline of the paper is as follows. In Sec. II, we derive the AdS/CFT dictionary and the laws for the CFT residing in the boundary dual to the five-dimensional Gauss-Bonnet AdS black hole. In Sec. III, we investigate the phase behavior and criticality of the CFT. Finally, in the last section we discuss the results and conclude them.
II Holographic thermodynamics of five-dimensional Gauss-Bonnet AdS black hole
The thermodynamic for black holes in AdS space-time has attracted a lot of attention and also raise perplexities. There are diverse perspectives on the thermodynamics of five-dimensional Gauss-Bonnet AdS black holes. From the viewpoint that the Gauss-Bonnet coupling constant is treated as an independent thermodynamic variable, the thermodynamics and phase transitions of Gauss-Bonnet AdS black holes have been systematically studied [41, 42, 43]. Adopting the perspective that Newton’s constant is allowed to vary and introducing a central charge, Kumar et al. derived a first law of thermodynamics that includes contributions from both bulk and boundary variables, known as the mixed form of the first law [44]. By varying Newton’s constant further and proposing the existence of two central charges: $A$-charge and $C$-charge, Fan et al. developed a new first law of black hole thermodynamics for Gauss-Bonnet AdS black holes [45].
In this section, we explore the extended thermodynamics of the five-dimensional neutral Gauss-Bonnet AdS black hole, based on the perspective that the cosmological constant and the Gauss-Bonnet coupling constant are treated as independent thermodynamic variables, while Newton’s constant is fixed [41, 42], we tentatively establish a connection between the extended thermodynamics of the five-dimensional neutral Gauss-Bonnet AdS black hole and the corresponding extended thermodynamics of dual CFT. By exploring the relationship between these two realms, we aim to shed light on the profound similarities and underlying principles that govern their thermodynamic properties.
II.1 Extended bulk thermodynamics
In this subsection, we briefly review the main points of the extended black hole thermodynamics for the five-dimensional neutral Gauss-Bonnet AdS black hole. A $d$-dimensional neutral Gauss-Bonnet AdS black hole is a solution for the gravitational field equation derived by the action [46]
@@ II.1 Extended bulk thermodynamics
Within the context of the low energy effective action formalism in heterotic string theory, the Gauss–Bonnet term is the leading order quantum correction to gravity [47]. The Gauss–Bonnet coupling constant is related to the string scale and can be identified with the inverse of the string tension [48, 41, 49]. Hence, we only consider positive Gauss-Bonnet coupling constant in this paper.
The theory admits an exact neutral black hole solution in five dimensions. The metric for the five-dimensional neutral Gauss-Bonnet AdS black hole is [50]
$ds^{2}=-f(r)dt^{2}+f^{-1}(r)dr^{2}+r^{2}d\Omega_{3}^{2},$ (5)
and the metric function $f(r)$ is given by
@@
From the viewpoint of the extended black hole thermodynamics, the cosmological constant $\Lambda$ is interpreted as the thermodynamic pressure and the black hole mass is regarded as the thermodynamic enthalpy $H$ [11]. The thermodynamic pressure is
$P=-\frac{\Lambda}{8\pi G_{\text{N}}}=\frac{3}{4\pi G_{\text{N}}}\frac{1}{L^{2}}.$ (10)
From the perspective of extended black hole thermodynamics, the thermodynamic volume can be defined either as a surface integral of the Killing potential [11] or as the Killing volume [51, 39]
[…]
where $\mid\mid\xi\mid\mid=\sqrt{-\xi\cdot\xi}$ is the norm of the event horizon generating Killing vector $\xi=\partial_{t}$. The thermodynamic volume for the five-dimensional Gauss-Bonnet AdS black hole is
@@
In the low-energy effective action of heterotic string theory, the Gauss-Bonnet coupling constant can be identified with the inverse of the string tension, making it natural to treat the Gauss-Bonnet constant as a thermodynamic variable. Furthermore, in order to obtain a consistent first law of black hole thermodynamics and the Smarr relation for the Gauss-Bonnet AdS black hole while the negative cosmological constant is regarded as the thermodynamic pressure and Newton’s constant is treated as a fixed parameter, the Gauss-Bonnet coupling constant $\alpha$ must be treated as a thermodynamic variable. The corresponding conjugate thermodynamic variable is given by
@@
The result is consistent with the Euclidean path integral approach, as presented in [52, 53], and is also in agreement with the Wald entropy formulation [54]. Evidently, the presence of the Gauss-Bonnet correction to Einstein’s theory of gravity leads to a notable revision of the Bekenstein-Hawking area law when considering the black hole’s temperature as its surface gravity, as emphasized in Ref. [55].
In the perspective of the fixed Newton’s constant, the consistent first law of black hole thermodynamics and the Smarr relation for the five-dimensional Gauss-Bonnet AdS black hole are [42]
@@
Unlike the perspective where Newton’s constant is treated as a variable, the first law and Smarr relation in this framework are independent of the boundary variables. Furthermore, treating the Gauss-Bonnet coupling constant as an independent thermodynamic variable is consistent with recent derivations using the extended Iyer-Ward formalism, where the coupling constant is interpreted as a conserved charge in black hole thermodynamics [22].
The first law and the Smarr relation are the starting point of our derivation of the first law of boundary thermodynamics. Based on the first law of black hole thermodynamics and the Smarr relation, we proceed to derive the boundary thermodynamics in the next subsection. This derivation establishes a tentative relationship between the thermodynamic properties of the bulk system and its boundary counterpart, allowing for a comprehensive analysis of their interconnected behavior. The insights gained from this derivation might enhance our understanding of the overall thermodynamic framework.
II.2 Extended boundary thermodynamics
Recently, Ahmed et al. proposed an approach that treats the conformal factor as a freely varying parameter, to solve the problem of the fixed Newton’s constant causing interdependence between the central charge and boundary volume. This perspective yields a holographic first law that corresponds to the extended thermodynamics for charged rotating AdS black holes [36]. In this subsection, we extend the methodology developed in [36] to the Gauss-Bonnet AdS black hole, deriving a holographic first law dual to Eq. (15).
From the perspective of the AdS/CFT correspondence, the dual CFT lives on the conformal boundary of the asymptotically AdS spacetime. According to the recent development of the holographic chemistry [24], the boundary metric of the CFT dual to the five-dimensional Gauss-Bonnet AdS spacetime is [2, 26]:
@@ II.2 Extended boundary thermodynamics
where $\omega$ is a dimensionless conformal factor. The standard approach is to set $\omega=1$, in which case the boundary volume $\mathcal{V}$ is proportional to $L^{3}$. Consequently, variations in the cosmological constant imply variations in the boundary CFT volume $\mathcal{V}$, suggesting the presence of a pressure-volume work term $\mathcal{P}d\mathcal{V}$, in the first law of CFT thermodynamics. However, this leads to the conclusion that the central charge $C$ and the thermodynamic volume $\mathcal{V}$ are not truly independent [36]. As proposed in [36], we instead choose the conformal factor $\omega=R/L$, which is allowed to vary freely [30, 36, 39], where $R$ is the radius of the boundary. This makes the boundary volume and the central charge independent thermodynamic variables. The variation of the conformal factor reflects the conformal symmetry of the boundary theory.
From the first law of black hole thermodynamics (15) and the Smarr relation (16) for the five-dimensional Gauss-Bonnet AdS black hole, we can express the first law in terms of the boundary thermodynamics
$\delta\left(\frac{M}{\omega}\right)=\left(\frac{T}{\omega}\right)\delta S+\frac{M-TS-\mathcal{A}\alpha}{\omega}\frac{\delta\left(L^{3}/G_{\text{N}}\right)}{\left(L^{3}/G_{\text{N}}\right)}\\ +\frac{\sqrt{G_{\text{N}}}\mathcal{A}}{\omega L}\delta\left(\frac{\alpha L}{\sqrt{G_{\text{N}}}}\right)-\left(\frac{\tilde{E}}{3\mathcal{V}}\right)\delta\mathcal{V},$ (18)
where $\mathcal{V}=2\pi^{2}R^{3}$ is the spatial volume of the boundary on which the CFT resides. In the derivation of the first law in terms of the boundary thermodynamics, the Newtown constant $G_{\text{N}}$ is regarded as a fixed constant, but the conformal factor $\omega$ is free to vary.
In the gauge/gravity duality, the central charge of the CFT is related to the AdS radius $L$ and the gravitational coupling constant $G_{\text{N}}$ in the bulk theory. As the authors of Ref. [44] did for the mixed first law of black hole thermodynamics for Gauss-Bonnet AdS black hole, the central charge is chosen as
$C=\frac{\Omega_{d-2}L^{d-2}}{16\pi G_{\text{N}}},$ (19)
where $\Omega_{d-2}$ is the volume of a $(d-2)$-dimensional unit sphere. Since the precise coefficient of the central charge is irrelevant for our purpose and only the combination $dC/C$ appears in the first law of the CFT, the dimensional dependent coefficient is normalized such that it agrees with the coefficient $A$ of the Euler density in the trace anomaly [28, 56, 57].
To get a consistent formulation of the first law of the CFT, we employ the following holographic dictionary:
$\begin{split}\tilde{E}&=\frac{M}{\omega},\qquad\qquad\tilde{T}=\frac{T}{\omega},\qquad S=S,\\ \tilde{\mathcal{A}}&=\frac{\sqrt{G_{\text{N}}}\mathcal{A}}{\omega L},\qquad\tilde{\mathcal{\alpha}}=\frac{\alpha L}{\sqrt{G_{\text{N}}}},\quad\mathcal{P}=\frac{\tilde{E}}{3\mathcal{V}},\\ C&=\frac{\pi L^{3}}{8G_{\text{N}}},\qquad\mu=\frac{1}{C}\left(\tilde{E}-\tilde{T}S-\tilde{\mathcal{A}}\tilde{\mathcal{\alpha}}\right),\end{split}$ (20)
@@
    $\begin{split}\tilde{\alpha}&=\sqrt{\frac{8CL^{3}}{\pi}}y,\\ \tilde{\mathcal{A}}&=-\frac{1}{R}\sqrt{\frac{\pi C}{8L^{3}}}\frac{24x^{4}+9x^{2}-6y}{x^{2}+2y}.\end{split}$ (29)
@@
III Thermodynamic ensembles in the dual CFT
The extended thermodynamics and phase transition for CFT states that are dual to the singly-spinning asymptotically AdS black hole were investigated by Ahmed et al., and various phase transitions and critical behavior in the CFT were uncovered [39]. Gong et al. found that the critical exponents for the CFT states in the vicinity of the critical point are the same with that of the Kerr-Newman AdS black hole [38]. In this section, we investigate the phase behavior and criticality exhibited by the canonical ensemble within the CFT framework.
III.1 Canonical ensemble: swallowtail behavior and critical point
The canonical ensemble for five-dimensional Gauss-Bonnet AdS black hole has been well study in the framework of extended black hole thermodynamics, i.e. black hole chemistry [42, 58]. In these studies it was found that the five-dimensional Gauss-Bonnet AdS black hole displays the swallowtail behavior, which indicates the small/large black hole phase transition reminiscent of the Van der Waals like phase transition. Besides, it was shown that the critical point for the five-dimensional Gauss-Bonnet AdS black hole phase transition is [42]